# Edge Universality for Inhomogeneous Random Matrices

### v1

@@ 1.1 Inhomogeneous random matrices
Random Matrix Theory (RMT), which lies at the intersection of matrix theory and probability theory—both ubiquitous in mathematics, science, and engineering—plays a central role in numerous areas of mathematics and physics; see the excellent survey [ABDF11]. When regarded as non-commutative random variables, the eigenvalues of random matrices form strongly correlated and highly complex systems. Remarkably, this correlation structure exhibits universality, depending only on the symmetry type of the matrix. This universality class, commonly referred to as random matrix statistics, differs significantly from the Gaussian universality class of the classical Central Limit Theorem [Dei07]. Such universality phenomena lie at the very core of RMT and extend far beyond the realm of random matrices themselves, including the KPZ universality class [Cor12, KPZ86] in stochastic surface growth and the BGS conjecture in quantum chaos [BGS84]. Now, random matrices have proven to be invaluable tools, with critical applications in number theory, statistics, machine learning, population dynamics, wireless communications, and mathematical finance [AGZ10, BS10, EY17, For10].
The physical origins of RMT can be traced back to Wigner’s seminal work in the 1950s [Wig55], in which he proposed large random Hamiltonians—including both mean-field and band matrices with random signs (Rademacher entries)—as models for complex quantum systems. Later, Wigner refined two fundamental questions concerning random matrices [Wig67]:
@@
While universality is well understood for mean-field ensembles—such as Wigner matrices [EY17], with nearly identical entry variances, and invariant ensembles [DG09], with conjugation-invariant law—real-world systems typically deviate from these idealized settings. Prominent examples—such as random band matrices [Bou18], adjacency matrices of sparse networks [LLV18], and structured covariance matrices [EK18, Joh07]—exhibit strong inhomogeneity and locality, which present substantial analytical challenges. This naturally raises a question: why can mean-field ensembles still effectively predict the behavior of such non-mean-field systems?
@@
Recently, inhomogeneous random matrices have attracted considerable attention, with notable progress on spectral norms [LvHY18, vH17] and non-asymptotic analysis [BCSvH24, BBvH23, BvH24a, BvH24b]. Despite these advances, the universality of local eigenvalue statistics—crucial for both theoretical insights and practical applications—remains far less understood in inhomogeneous settings.
@@ 1.2 Model and main results
Let $W_{N}$ be an $N\times N$ real symmetric (or complex Hermitian) Wigner matrix, let $\Sigma_{N}$ be a deterministic variance-profile matrix, and let $A_{N}$ be a finite-rank perturbation. Define
@@
    (Variance profile) The profile matrix $\Sigma_{N}=(\sigma_{ij})$ has nonnegative entries, and its squared entries form a stochastic matrix $\sum_{j=1}^{N}\sigma_{ij}^{2}=1$ for each $i$. Equivalently, $P_{N}:=(\sigma_{ij}^{2})_{i,j=1}^{N}$ is a Markov transition matrix.
@@
To establish the universality of Tracy-Widom distributions [TW94, TW96] and their BBP deformations [BBAP05, BV13, BV16], we introduce a key concept: the mixing property of a Markov chain $([N],P_{N})$, which characterizes its convergence rate to the uniform distribution. We remark that $[N]$ can be replaced throughout by an arbitrary finite state space $S$ (with $N=|S|$), and the index $i\in[N]$ by any $i\in S$.
@@
    (Long-time uniform mixing) There exists $\delta\in(0,0.1)$, such that for all $n\geq t_{N}$, and all $x,y\in[N]$,
@@
Our principal theorem establishes the Tracy-Widom law and BBP transition for a broad class of deformed inhomogeneous random matrices.
@@
Conditions (B1) and (B2) are both essential for edge universality, and the violation of either leads to non-universal behavior. For instance, the block-diagonal model satisfies Condition (B1) but violates Condition (B2), resulting in a clear failure of universality at the edge. Conversely, in dimensions $d>2$, random band matrices satisfy Condition (B2) but violate Condition (B1), necessitating a non-universal correction at the spectral edge, as shown in [LZ24b].
Furthermore, the moment condition $\theta t_{N}\ll N^{\frac{1}{3}}$ is also sharp. In the sparse Wigner case with $t_{N}=1$, if $\theta\gg N^{1/3}$, a non-universal correction emerges, as demonstrated in [LS18].
@@
Our universality results are consistent with the Thouless criterion for localization, adapted to the band matrix setting by Fyodorov and Mirlin [FM94]; see also [Sod10, Spe11, Wan92]. Informally, the Thouless criterion [Tho77] compares a dynamical (diffusive/mixing) time scale with the spectral resolution scale and thereby predicts whether local spectral statistics follow GOE/GUE laws or degenerate to Poisson statistics.
More concretely, let $t_{\mathrm{mix}}:=t_{N}$ denote the mixing time of the variance-profile Markov chain $([N],P_{N})$ and define the Thouless time as
$t_{\mathrm{Th}}:=\frac{\varrho(\alpha_{0})}{\Delta(\alpha_{0})}.$ (1.10)
One expects the following dichotomy:
@@
Our long-time mixing condition (B2) in Theorem 1.3 ensures the model lies firmly in the regime where $t_{\mathrm{mix}}\ll t_{\mathrm{Th}}$. For further physical intuition regarding Conditions (B1) and (B2), we refer to [GMGW98, Section VI, (iii)].
@@ 1.3 Previous results
Local universality for Wigner matrices is by now well-established, with foundational breakthroughs achieved over the past three decades [EPR+10, ESY11, Joh01, Sos99, TV10, TV11], followed by extensions to generalized Wigner matrices [BEYY16, EYY12, BEY14]. We refer to surveys such as [AGZ10, Bou18, EY17] for comprehensive overviews. Parallel advances have established universality for sparse random matrices: in the bulk spectrum [EKYY13b, HLY15], and at the spectral edge after appropriate corrections [EKYY12, HLY20, HK21, HY22, LS18]. For random $d$-regular graphs, bulk universality is known for $d\geq N^{\epsilon}$ [BHKY17], and edge universality for all $d\geq 3$ was recently proved in [HMY24]. Later, universality and delocalization on some specifics non-mean-field random matrix models, typically random band matrices, were explored. Important contributions include [BE17, BYY20, BGP14, EK11b, EK11a, EKYY13a, HM19, LZ24b, Sod10, SS21, Shc14, YYY21, YYY22, XYYY24] (though this list is not exhaustive). Very recent breakthrough results on bulk universality and delocalization of random band matrices have been established in [DYYY25b, DYYY25a, ER25, FYY25, LZ24b, XYYY24, YY25b, YY25c]. Most focus on the specific block variance profile or complex Gaussian entries, with the exception of the one-dimensional case [ER25]. We refer to [ER25, DYYY25b] and references therein for state-of-the-art results and surveys.
Edge universality serves as the central focus of this paper. The study of spectral edge statistics originated in the seminal contributions of Tracy and Widom [TW94, TW96], Forrester [For93], which introduced the now ubiquitous Tracy-Widom distributions and Airy point processes for GUE and GOE ensembles. These limiting distributions were later shown to hold for general Wigner matrices under suitable moment conditions [Sos99, Joh12, TV10], culminating in a necessary and sufficient characterization established by Lee and Yin [LY14]. A parallel line of inquiry concerns universality in deformed random matrices and the BBP phase transition [BBAP05, Péc06, BV13, BV16, LS15]; we refer to [Péc14] for a comprehensive survey.
@@
Inhomogeneous (or structured) random matrices arise naturally in a wide range of applications, such as statistical inference, numerical linear algebra, compressed sensing, and network analysis. These have attracted significant interest in studies of extreme singular values [DZ24, LTV21, RV10], spectral outliers [GLZ24, Han24b, TY21], operator norms [BvH16, LvHY18, vH17], and concentration or deviation inequalities [BBvH23, BCSvH24, Tro15]. For comprehensive overviews, we refer to [BvH24a, Tro15, Ver12, vH17].
For Wigner-type inhomogeneous matrices, a growing body of research has addressed both global and extremal spectral properties. In particular, work in [Chi23, GNT15] demonstrated that the empirical spectral distribution converges to the semicircle law under the mild condition that $\max_{x,y}\sigma_{xy}\to 0$. Under sharp moment conditions, almost sure convergence of the largest eigenvalue was established in [ASTY24]. The behavior of outliers in deformed inhomogeneous ensembles has been investigated in [GLZ24, Han24b], and several non-asymptotic deviation bounds for the operator norm have been derived in [BvH24a].
Although much of the literature has focused on non-asymptotic bounds, such results are often suboptimal and may fail to capture the correct fluctuation scale in general settings (see [BvH24a, Section 1.2.2]). In contrast, asymptotic fluctuation theory—such as Tracy-Widom-type laws for the spectral edge—remains underdeveloped, with the exception of results on outlier fluctuations in [GLZ24].
@@
In essence, while the method of moments is central to our approach, it requires the analysis of very high-order moments, which presents significant technical challenges. The method of moments in probability theory was first rigorously treated by P. L. Chebyshev in 1887 [Tch90], who used it to provide the first fully rigorous proof of the central limit theorem [Dia87, Bil12]. In random matrix theory, Wigner pioneered this method in 1955 to establish the semicircle law [Wig55, Wig58], and it was later employed by Soshnikov in 1999 to prove the Tracy-Widom law [Sos99]. Subsequently, Feldheim and Sodin [FS10] developed a more sophisticated Chebyshev polynomial moment approach, extending universality results to both Wigner and sample covariance matrices. In a related direction, Okounkov [Oko00] rederived the Tracy-Widom law and the corresponding Laplace transform for the GUE ensemble while studying the longest increasing subsequence problem in random permutations.
A major breakthrough was achieved in Sodin’s seminal work [Sod10], which identified the critical bandwidth scaling $W_{c}=L^{5/6}$ for one-dimensional random band matrices with unimodular entries and a cutoff variance profile. This work rigorously characterized edge statistics in both extended and localized regimes. Subsequently, the Chebyshev polynomial framework was extended into a polynomial moment method by [LZ24b], enabling the analysis of edge statistics for random band matrices in higher dimensions. Furthermore, the method of large moments has also been applied to investigate outliers of deformed IRM ensembles in very general settings [GLZ24].
@@ 1.4 Key contributions
This paper highlights the powerful polynomial moment method, developing a general strategy to resolves previously inaccessible universality problems for random matrices with arbitrary variance profiles and arbitrary entry distributions, requiring only independence of the entries under the essential sole assumption of independence. This approach builds upon the combinatorial and diagrammatic expansions techniques developed in [FS10, Sod10, EK11a, LZ24b, GLZ24], employing a Chebyshev polynomial moment framework.
@@
    New Chebyshev expansion for inhomogeneous Gaussian matrices. We establish a closed-form diagrammatic expansion for Chebyshev polynomial moments in inhomogeneous Gaussian ensembles by combining Wick’s formula with classical identity for Chebyshev polynomials (see Proposition 2.5 and Theorem 2.6 below). Unlike previous methods that rely on combinatorial recursions and non-backtracking paths within the unimodular framework [FS10, Sod10, EK11a, LZ24b], our approach reveals a new combinatorial identity unique to the Gaussian setting. This formulation provides a clearer and more efficient reduction from trace moments to diagrammatic (i.e., reduced ribbon graph) representations in the Gaussian case. Whereas diagram functions in the unimodular framework are only introduced after an approximation step, as in [Sod10, LZ24b], our Gaussian-specific expansion avoids such intermediate approximations.
@@
    Wide-ranging applicability and sharpness. Our universality results substantially extend recent important advances on random band matrices in the delocalized regime—previously limited to unimodular entries and highly structured variance profiles [Sod10, LZ24b]. Moreover, our framework encompasses a broad spectrum of random matrix ensembles, including highly inhomogeneous, sparse, and non-mean-field models, extending well beyond classical settings. As detailed in Section 7, Theorem 1.3 in several cases establishes sharp criteria for edge universality.
@@ 1.5 Structure of this paper
Notation. Throughout this paper, constants such as $C,C_{1},\ldots$ are independent of $N$. $f(n)=o(g(n))$ or $f(n)\ll g(n)$ means that ${f(n)}/{g(n)}\to 0$ as $n\rightarrow\infty$, while $f(n)=O(g(n))$ means that f(n)/g(n){f(n)}/{g(n)} is bounded. $a\wedge b=\min\{a,b\}$, $a\lor b=\max\{a,b\}$.

### v2

@@ 1.1 Inhomogeneous random matrices
Random Matrix Theory (RMT), which lies at the intersection of matrix theory and probability theory—both ubiquitous in mathematics, science, and engineering—plays a central role in numerous areas of mathematics and physics; see the excellent survey [akemann2011oxford]. When regarded as non-commutative random variables, the eigenvalues of random matrices form strongly correlated and highly complex systems. Remarkably, this correlation structure exhibits universality, depending only on the symmetry type of the matrix. This universality class, commonly referred to as random matrix statistics, differs significantly from the Gaussian universality class of the classical Central Limit Theorem [MR2334189]. Such universality phenomena lie at the very core of RMT and extend far beyond the realm of random matrices themselves, including the KPZ universality class [MR2930377, PhysRevLett.56.889] in stochastic surface growth and the BGS conjecture in quantum chaos [BGS1984]. Now, random matrices have proven to be invaluable tools, with critical applications in number theory, statistics, machine learning, population dynamics, wireless communications, and mathematical finance [MR2760897, bai2010spectral, erdos2017dynamical, forrester2010log].
The physical origins of RMT can be traced back to Wigner’s seminal work in the 1950s [Wigner1955Characteristic], in which he proposed large random Hamiltonians—including both mean-field and band matrices with random signs (Rademacher entries)—as models for complex quantum systems. Later, Wigner refined two fundamental questions concerning random matrices [wigner1967random]:
@@
While universality is well understood for mean-field ensembles—such as Wigner matrices [erdos2017dynamical], with nearly identical entry variances, and invariant ensembles [MR2514781], with conjugation-invariant law—real-world systems typically deviate from these idealized settings. Prominent examples—such as random band matrices [bourgade2018random], adjacency matrices of sparse networks [MR3966517], and structured covariance matrices [MR3966514, MR2334195]—exhibit strong inhomogeneity and locality, which present substantial analytical challenges. This naturally raises a question: why can mean-field ensembles still effectively predict the behavior of such non-mean-field systems?
@@
Recently, inhomogeneous random matrices have attracted considerable attention, with notable progress on spectral norms [MR3878726, MR3837269] and non-asymptotic analysis [bandeira2024matrix, MR4635836, brailovskaya2024extremal, MR4823211]. Despite these advances, the universality of local eigenvalue statistics—crucial for both theoretical insights and practical applications—remains far less understood in inhomogeneous settings.
@@ 1.2 Model and main results
Let $W_{N}$ be an $N\times N$ real symmetric (or complex Hermitian) Wigner matrix, let $\Sigma_{N}$ be a deterministic variance-profile matrix and $A_{N}$ be a finite-rank perturbation. Define
@@
    (Variance profile) The profile matrix $\Sigma_{N}=(\sigma_{ij})$ has nonnegative entries such that $\sum_{j=1}^{N}\sigma_{ij}^{2}=1$ for each $i$. Equivalently, $P_{N}:=(\sigma_{ij}^{2})_{i,j=1}^{N}$ is a Markov transition matrix.
@@
To establish the universality of Tracy-Widom distributions [TW1994, TW1996] and their BBP deformations [baik2005phase, bloemendal2013limits, bloemendal2016limits], we introduce a key concept: the mixing property of a Markov chain $([N],P_{N})$, which characterizes its convergence rate to the uniform distribution on $[N]$. We remark that $[N]$ can be replaced throughout by an arbitrary finite state space $S$ (with $N=|S|$), and the index $i\in[N]$ by any $i\in S$.
@@
    (Long-time uniform mixing) There exists $\delta\in(0,0.1)$, such that for all $n\geq t_{N}$ and all $x,y\in[N]$,
@@
Our principal theorem establishes the Tracy-Widom law and BBP transition for a broad class of deformed inhomogeneous random matrices. This can be interpreted as a Lindeberg-type central limit theorem for the Tracy-Widom law within random matrix theory, even in the non-deformed case.
@@
Conditions (B1) and (B2) are both essential for edge universality, and the violation of either leads to non-universal behavior. For instance, the block-diagonal GUE model satisfies Condition (B1) but violates Condition (B2), resulting in a clear failure of universality at the edge. Conversely, in dimensions $d>2$, random band matrices satisfy Condition (B2) but violate Condition (B1), necessitating a non-universal correction at the spectral edge, as shown in [liu2023edge].
Furthermore, the condition $\theta t_{N}\ll N^{{1}/{3}}$ is also sharp. In the sparse Wigner case with $t_{N}=1$, if $\theta\gg N^{1/3}$, a non-universal correction emerges, as demonstrated in [lee2018].
@@
Our universality results are consistent with the Thouless criterion for localization, adapted to the band matrix setting by Fyodorov and Mirlin [fyodorov1994statistical]; see also [sodin2010spectral, spencer2011random, Wang1992]. Informally, the Thouless criterion [thouless1977maximum] compares a dynamical (diffusive/mixing) time scale with the spectral resolution scale and thereby predicts whether local spectral statistics follow GOE/GUE laws or degenerate to Poisson statistics.
More concretely, let $t_{\mathrm{mix}}:=t_{N}$ denote the mixing time of the variance-profile Markov chain $([N],P_{N})$ and let $t_{\mathrm{Th}}$ be the Thouless time. One expects the following dichotomy:
@@
Our long-time mixing condition (B2) in Theorem 1.3 ensures that the model lies firmly in the regime where $t_{\mathrm{mix}}\ll t_{\mathrm{Th}}$. For further physical intuition regarding Conditions (B1) and (B2), we refer to [guhr1998random, Section VI, (iii)].
@@ 1.3 Previous results
Local universality for Wigner matrices is by now well-established, with foundational breakthroughs achieved over the past three decades [erdHos2010bulk, erdHos2011universality, johansson2001universality, soshnikov1999universality, tao2010random, tao2011random], followed by extensions to generalized Wigner matrices [bourgade2016fixed, erdos2012generalized, MR3253704]. We refer to surveys such as [MR2760897, bourgade2018random, erdos2017dynamical] for comprehensive overviews. Parallel advances have established universality for sparse random matrices: in the bulk spectrum [MR3098073, MR3429490], and at the spectral edge after appropriate corrections [MR2964770, MR4089498, MR4288336, huang2022edge, lee2018]. For random $d$-regular graphs, bulk universality is known for $d\geq N^{\epsilon}$ [MR3729611], and edge universality for all $d\geq 3$ was recently proved in [huang2024ramanujan]. Later, universality and delocalization on some specific non-mean-field random matrix models, typically random band matrices, were explored. Important contributions include [bao2017delocalization, bourgade2020random, benaych2014largest, erdHos2011quantum, EK11Quantum, erdHos2013delocalization, he2019diffusion, liu2023edge, sodin2010spectral, Shcherbina2021, shcherbina2014second, yang2021delocalization, yang2022delocalization, MR4736267] (though this list is not exhaustive). Very recent breakthrough results on universality and delocalization of random band matrices have been established in [dubova2025delocalization, dubova2025delocalization2, erdHos2025zigzag, fan2025blockreductionmethodrandom, liu2023edge, MR4736267, yang2025delocalization2, yau2025delocalization]. Most focus on the specific block variance profile or complex Gaussian entries, with the exception of the one-dimensional case [erdHos2025zigzag]. We refer to [erdHos2025zigzag, dubova2025delocalization] and references therein for state-of-the-art results and surveys.
Edge universality serves as the central focus of this paper. The study of spectral edge statistics originated in the seminal contributions of Tracy and Widom [TW1994, TW1996], Forrester [forrester1993spectrum], which introduced the now ubiquitous Tracy-Widom distributions and Airy point processes for GUE and GOE ensembles. These limiting distributions were later shown to hold for general Wigner matrices under suitable moment conditions [soshnikov1999universality, johansson2012universality, tao2010random], culminating in a necessary and sufficient characterization established by Lee and Yin [MR3161313]. A parallel line of inquiry concerns universality in deformed random matrices and the BBP phase transition [baik2005phase, peche2006largest, bloemendal2013limits, bloemendal2016limits, lee2015edge]; we refer to [peche2014deformed] for a comprehensive survey.
@@
Inhomogeneous (or structured) random matrices arise naturally in a wide range of applications, such as statistical inference, numerical linear algebra, compressed sensing, and network analysis. These have attracted significant interest in studies of extreme singular values [MR4779853, MR4255145, MR2827856], spectral outliers [geng2024outliers, han2024outliers, MR4234995], operator norms [bandeira2016sharp, MR3878726, MR3837269], and concentration or deviation inequalities [MR4635836, bandeira2024matrix, tropp2015introduction]. For comprehensive overviews, we refer to [brailovskaya2024extremal, tropp2015introduction, Vershynin2012, MR3837269].
For Wigner-type inhomogeneous matrices, a growing body of research has addressed both global and extremal spectral properties. In particular, [MR4552698, MR3416062] demonstrated that the empirical spectral distribution converges to the semicircle law under the mild condition that $\max_{x,y}\sigma_{xy}\to 0$. Under sharp moment conditions, almost sure convergence of the largest eigenvalue was established in [altschuler2024spectral]. The behavior of outliers in deformed inhomogeneous ensembles has been investigated in [geng2024outliers, bandeira2024matrix], and several non-asymptotic deviation bounds for the operator norm have been derived in [brailovskaya2024extremal].
Although much of the literature has focused on non-asymptotic bounds, such results are often suboptimal and may fail to capture the correct fluctuation scale in general settings (see [brailovskaya2024extremal, Section 1.2.2]). In contrast, asymptotic fluctuation theory—such as Tracy-Widom-type laws for the spectral edge—remains underdeveloped, with the exception of results on outlier fluctuations in [geng2024outliers].
@@
In essence, while the method of moments is central to our approach, it requires the analysis of very high-order moments, which presents significant technical challenges. The method of moments in probability theory was first rigorously treated by P. L. Chebyshev in 1887 [chebyshev1890deux], who used it to provide the first fully rigorous proof of the central limit theorem [diaconis11987application, billingsley]. In random matrix theory, Wigner pioneered this method in 1955 to establish the semicircle law [Wigner1955Characteristic, wigner1958distribution], and it was later employed by Soshnikov in 1999 to prove the Tracy-Widom law [soshnikov1999universality]. Subsequently, Feldheim and Sodin [feldheim2010universality] developed a more sophisticated Chebyshev polynomial moment approach, extending universality results to both Wigner and sample covariance matrices. While studying the longest increasing subsequence problem in random permutations, Okounkov [okounkov2000random] rederived the Tracy-Widom law and the corresponding Laplace transform for the classical GUE ensemble.
A major breakthrough was achieved in Sodin’s seminal work [sodin2010spectral], which identified the critical bandwidth scaling $W_{c}=L^{5/6}$ for one-dimensional random band matrices with unimodular entries and a cutoff variance profile. This work rigorously characterized edge statistics in both extended and localized regimes. Subsequently, the Chebyshev polynomial framework was extended into a polynomial moment method by [liu2023edge], enabling the analysis of edge statistics for random band matrices in higher dimensions. Furthermore, the method of large moments has also been applied to investigate outliers of deformed IRM ensembles in very general settings [geng2024outliers].
@@ 1.4 Key contributions
This paper highlights the powerful polynomial moment method, developing a general strategy to resolve previously inaccessible universality problems for random matrices with arbitrary variance profiles and arbitrary entry distributions, requiring only the essential sole assumption of independence. This approach builds upon the combinatorial and diagrammatic expansion techniques developed in [feldheim2010universality, sodin2010spectral, EK11Quantum, liu2023edge, geng2024outliers], employing a Chebyshev polynomial moment framework.
@@
    New Chebyshev expansion for inhomogeneous Gaussian matrices. We establish a closed-form diagrammatic expansion for Chebyshev polynomial moments in inhomogeneous Gaussian ensembles by combining Wick’s formula with classical identity for Chebyshev polynomials (see Proposition 2.5 and Theorem 2.6 below). Unlike previous methods that rely on combinatorial recursions and non-backtracking paths within the unimodular framework [feldheim2010universality, sodin2010spectral, EK11Quantum, liu2023edge], our approach reveals a new combinatorial identity unique to the Gaussian setting. This formulation provides a clearer and more efficient reduction from trace moments to diagrammatic (i.e., reduced ribbon graph) representations in the Gaussian case. Whereas diagram functions in the unimodular framework are only introduced after an approximation step, as in [sodin2010spectral, liu2023edge], our Gaussian-specific expansion avoids such intermediate approximations.
@@
    Wide-ranging applicability and sharpness. Our universality results substantially extend recent important advances on random band matrices in the delocalized regime—previously limited to unimodular entries and highly structured variance profiles [sodin2010spectral, liu2023edge]. Moreover, our framework encompasses a broad spectrum of random matrix ensembles, including highly inhomogeneous, sparse, and non-mean-field models, extending well beyond classical settings. As detailed in Section 7, Theorem 1.3 in several cases establishes sharp criteria for edge universality.
@@ 1.5 Structure of this paper
Notation. Throughout this paper, constants such as $C,C_{1},\ldots$ are independent of $N$. $f(n)=o(g(n))$ or $f(n)\ll g(n)$ means that ${f(n)}/{g(n)}\to 0$ as $n\rightarrow\infty$, while $f(n)=O(g(n))$ means that f(n)/g(n){f(n)}/{g(n)} is bounded. The symbols $a\wedge b=\min\{a,b\}$, $a\lor b=\max\{a,b\}$.